# On the verification of a Nicolas inequality

### v2

@@ Abstract.
where $x\geq 2$, $N_{x}$ denotes the product of the primes less or equal than $x$, $\gamma$ is the Euler constant and $\varphi$ is the Euler totient function. We see that verification of (1) depends on the sign of the Big-o function in the Mertens estimate for the sum of reciprocals of primes that. Then, we analyze the value of such an error term.
@@ 1. Introduction and notation
Nicolas inequality was introduced in [7, Theorem 2] and establish that one and only one of the following holds:
@@
Let us introduce some terminology. We have used standard notation for number theory that can be found, for instance, in [8, 11]. Given a positive function $g(x)$, the Big-o function of $g(x)$, denoted by $O(g(x))$, is a function such that $|O(g(x))|\leq Cg(x)$. Observe that necessarily $C>0$, but $O(g(x))$ could be negative, positive or both. Throughout the paper we consider $x$ as a positive real variable, $p$ denotes prime numbers and $n$ and $d$ denote natural numbers. Prime zeta function is defined as the sum of the powers of reciprocals of prime numbers, i.e. $P(s)=\sum_{p}\frac{1}{p^{s}}$, where $s$ is a complex number, (see for instance [2, 5]), also $P_{x}(s)=\sum_{p\leq x}\frac{1}{p^{s}}$. The von Mangoldt function $\Lambda(n)$ is the arithmetic function $\Lambda(n)=\log p$ if $n=p^{k}$ for some positive integer $k$, $\Lambda(n)=0$ otherwise. The so-called Chebyshev functions are $\psi(x)=\sum_{n\leq x}\Lambda(n)$ and $\theta(x)=\sum_{p\leq x}\log p$, and the prime counting function is $\pi(x)=\sum_{p\leq x}1$. We define $N_{x}=\prod_{p\leq x}p$ and $\varphi_{x}=\prod_{p\leq x}(p-1)$. Observe that $\theta(x)=\log N_{x}$ and $\varphi_{x}=\varphi(N_{x})$, where $\varphi(n)=|\{m\in\mathbb{N}:m\leq n,(m,n)=1\}|$ is the Euler totient function. The floor function is $x]=\max\{m\in\mathbb{Z}:m\leq x\}$, so we can write $x]=x-\{x\}$, where function $\{x\}$ is bounded between 0 and $1$.
@@ 2. On the sign of the error term
In this unique section we prove a theorem and discuss the error term. The first one relates the sign of the error term of Mertens theorem for the sum of reciprocal of primes with the verification of the Nicolas inequality. To this aim we apply results involving the zeta prime function, basic estimates for Chevyshev functions and standard techniques of real analysis. In the second part we analyze the sign of such an error term, to this end we revise the proof of Mertens Theorem and obtain an explicit formula for the term, then we approximate sufficiently its value, however we were not able to determine the sign.
@@
(II) If there are $x_{0}\geq 2$ and $C<0$ such that $O(1/\log x)<\frac{C}{\log x}$ for every $x\geq x_{0}$, then (F) holds.
@@
Taking into account Mertens Theorem (see e.g. [8, Theorem 8.8(d)]) and the above equation, we can rewrite the Nicolas inequality, namely $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi_{x}}$, as
@@
where $\gamma$ is the Euler constant and $b$ is known as the Mertens constant (see [1] and [10, (2.7)]).
@@
Therefore $x^{e^{C/\log x}}>x(1+\frac{\alpha}{\log x})$ and for a suitable $\alpha$ we can ensure that $x^{e^{C/\log x}}>\theta(x)$ for every $x\geq x_{2}$, we can choose for instance $\alpha=1/2$ (see [10, Theorem 4(3.15)]). In consequence, taking logarithms twice at both sides, we first get $e^{C/\log x}\log x>\log\theta(x)$, and then $\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}>0$ for every $x\geq x_{2}$. This implies that there exists some $\varepsilon>0$ such that
@@
Therefore Nicolas inequality holds for every $x\geq x_{3}$, which is not exactly (T), but it is in contradiction with (F), and as by [7, Theorem 2] there are only two possibilities, (T) must hold.
Step 3: (II). Now $C<0$ and by hypothesis there is $x_{0}$ such that $O(1/\log x)<\frac{C}{\log x}<0$ for every $x\geq x_{0}$, moreover we know there is some $x_{4}\geq x_{0}$ such that $\theta(x_{4})>x_{4}$ (see, for instance, the first page in [9]), join these facts with reasonings in previous steps, we obtain
@@ 2.1. Discussion on the sign of the error term
Step 1. Let start with an estimate for the sum of logarithms (see [8, Lemma 8.2]).
$\sum_{n\leq x}\log x=x\log x+x+O_{1}(\log x)\,$,
where
$O_{1}(\log x)=1-\Big(\{x\}\log x-\int_{1}^{x}\dfrac{\{t\}}{t}\,dt\Big)\,.$
So integration by parts brings to
$O_{1}(\log x)=1-\Big(\int_{1}^{x}\{t\}^{\prime}\log t\,dt\,\Big)\,.$
We see that $\{x\}=x-[x]$ is bounded, strictly increasing, derivable and integrable in $\mathbb{R}\setminus\mathbb{Z}$, thus $\{t\}^{\prime}$ is positive and bounded in $\mathbb{R}\setminus\mathbb{Z}$ and the integral of $\log(t)$ is an unbounded function. In consequence $O_{1}(\log x)$ is negative for sufficiently large $x$. More formally, there exists $x_{1}\geq 2$ such that $O_{1}(\log x)<0$ for all $x\geq x_{1}$.
Step 2. We follow with a sum involving the Mangoldt function (see [8, Theorem 8.8(a)]).
@@
for $k\in(10^{4},10^{5})$ and a list of primes $\{p_{j}\}_{j}$. Alternatively we found estimates from above and from below for the second term. We can assume that $x$ is integer, hence the reminder of $\dfrac{x}{d}$ is $1$ for $d\in(x/2,x)$, so $\Big\{\dfrac{x}{d}\Big\}=\dfrac{1}{d}$ for $d\in(x/2,x)$, also recall that $\Big\{\dfrac{x}{x}\Big\}=0$. Taking into account [8, Theorem 8.3], we can assert that $\psi(x/2)<3/4\psi(x)$ for a sufficient large integer $x$, thus
@@
In consequence, taking into account that $ax<\psi(x)<cx$ for $a=0.780355$... and $c=1.170533$..., for a sufficiently large $x$ (see [8, Theorem 8.3]), we deduce that
@@
Step 3. Next we deal with a sum over primes (see [8, Theorem 8.8(b)]).
@@
Let us obtain two coarse bounds. We know that the second term is bounded and, since $\log p/p^{n}$ is positive, it increases when the $x$ grows and the partial sums in $n$ grows. Moreover it is bounded from above by $\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}$ (see proof of [8, Theorem 8.8(b)]), which is bounded by the integral of $\dfrac{\log(x-1)}{(x-1)(x-2)}$ between $2$ and infinity. So, for sufficiently large $x$
@@
$1.645<O_{3}(1)<-0.048<0.585-0.633$
Step 4. Finally we study the error term of the sum of reciprocal of primes (see [8, Theorem 8.8(d)]). We define $R(x)=\displaystyle\sum_{p\leq x}\dfrac{\log p}{p}-\log x$, so thanks to the estimate given in previous steps we get $R(x)=O_{3}(1)$. Let
@@
$O(1/\log x)=\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\,.$

### v5

# On the “verification” of a Nicolas inequality

@@ Abstract.
where $x\geq 2$, $N_{x}$ denotes the product of the primes less or equal than $x$, $\gamma$ is the Euler constant and $\varphi$ is the Euler totient function. We see that verification of (1) depends on the sign of the big-O function in the Mertens estimate for the sum of reciprocals of primes. Then, we analyze the sign of such an error term.
@@
Note to the reader
First version of the present preprint has an error that brings us to an incorrect conclusion. In second version we amend this error, but we do not provide new contributions. In this new version we have corrected Step 1 of final discussion and added Remark 2.4, and also (as previously) we would advise the reader that the paper contains some imprecisions, that the author have left to the reader, and also could include some undetected mistakes, so the author appreciates any comment without attached files (in order to avoid fishing) to galdames@uv.es .
@@ 1. Introduction and notation
Nicolas inequality was introduced in [6, Theorem 2] and establish that one and only one of the following holds:
@@
Let us introduce some terminology. We have used standard notation for number theory that can be found, for instance, in [7, 10]. Given a positive function $g(x)$, the big-O function of $g(x)$, denoted by $O(g(x))$, is a function such that $|O(g(x))|\leq Cg(x)$ uniformly. Observe that necessarily $C>0$, otherwise $O(g(x))$ vanishes, but $O(g(x))$ could be negative, positive or both. Moreover we can write $O(g(x))=C(x)g(x)$ for some uniformly bounded function $C(x)$. Throughout the paper we consider $x$ as a positive real variable, $p$ denotes prime numbers and $n$ and $d$ denote natural numbers. Prime zeta function is defined as the sum of the powers of reciprocals of prime numbers, i.e. $P(s)=\sum_{p}\frac{1}{p^{s}}$, where $s$ is a complex number, (see for instance [1, 4]), also $P_{x}(s)=\sum_{p\leq x}\frac{1}{p^{s}}$. The von Mangoldt function $\Lambda(n)$ is the arithmetic function $\Lambda(n)=\log p$ if $n=p^{k}$ for some positive integer $k$, $\Lambda(n)=0$ otherwise. The so-called Chebyshev functions are $\psi(x)=\sum_{n\leq x}\Lambda(n)$ and $\theta(x)=\sum_{p\leq x}\log p$, and the prime counting function is $\pi(x)=\sum_{p\leq x}1$. We define $N_{x}=\prod_{p\leq x}p$ and $\varphi_{x}=\prod_{p\leq x}(p-1)$. Observe that $\theta(x)=\log N_{x}$ and $\varphi_{x}=\varphi(N_{x})$, where $\varphi(n)=|\{m\in\mathbb{N}:m\leq n,(m,n)=1\}|$ is the Euler totient function. The floor function is $x]=\max\{m\in\mathbb{Z}:m\leq x\}$, so we can write $x]=x-\{x\}$, where function $\{x\}$ is bounded between 0 and $1$.
@@ 2. On the sign of the error term
In this unique section we prove a theorem and discuss the error term. The first one relates the sign of the error term of Mertens theorem for the sum of reciprocal of primes with the verification of the Nicolas inequality. To this aim we apply results involving the zeta prime function, basic estimates for Chevyshev functions and standard techniques of real analysis. In the second part we analyze the sign of such an error term, to this end we revise the proof of Mertens Theorem and obtain an explicit formula for the term, then we approximate sufficiently its value.
@@
(II) If there are $x_{0}\geq 2$ and $C\leq 0$ such that $O(1/\log x)\leq\frac{C}{\log x}$ for every $x\geq x_{0}$, then (F) holds.
@@
Taking into account Mertens Theorem (see e.g. [7, Theorem 8.8(d)]) and the above equation, we can rewrite the Nicolas inequality, namely $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi_{x}}$, as
@@
where $\gamma$ is the Euler constant and $b$ is known as the Mertens constant ([9, (2.7)]).
@@
Therefore $x^{e^{C/\log x}}>x(1+\frac{\alpha}{\log x})$ and for a suitable $\alpha$ we can ensure that $x^{e^{C/\log x}}>\theta(x)$ for every $x\geq x_{2}$, we can choose for instance $\alpha=1/2$ (see [9, Theorem 4(3.15)]). In consequence, taking logarithms twice at both sides, we first get $e^{C/\log x}\log x>\log\theta(x)$, and then $\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}>0$ for every $x\geq x_{2}$. This implies that there exists some $\varepsilon>0$ such that
@@
Therefore Nicolas inequality holds for every $x\geq x_{3}$, which is not exactly (T), but it is in contradiction with (F), and as by [6, Theorem 2] there are only two possibilities, (T) must hold.
Step 3: (II). Now $C\leq 0$ and by hypothesis there is $x_{0}$ such that $O(1/\log x)<\frac{C}{\log x}\leq 0$ for every $x\geq x_{0}$, moreover we know there is some $x_{4}\geq x_{0}$ such that $\theta(x_{4})>x_{4}$ (see, for instance, the first page in [8]), join these facts with reasonings in previous steps, we obtain
@@ 2.1. Discussion on the sign of the error term
Step 1. Let start with an estimate for the sum of logarithms (see [7, Lemma 8.2 and Theorem 8.8(a)] and also [2]).
$\sum_{n\leq x}\log x=x\log x-x+O_{1}(\log x)\,$,
where do not mind the sign of $O_{1}(\log x)$ because in next step we will divide it by $x$.
Step 2. We follow with a sum involving the Mangoldt function (see [7, Theorem 8.8(a)]).
@@
for $k\in(10^{4},10^{5})$ and a list of primes $\{p_{j}\}_{j}$. Alternatively we found estimates from above and from below for the second term. We can assume that $x$ is integer, hence the reminder of $\dfrac{x}{d}$ is $1$ for $d\in(x/2,x)$, so $\Big\{\dfrac{x}{d}\Big\}=\dfrac{1}{d}$ for $d\in(x/2,x)$, also recall that $\Big\{\dfrac{x}{x}\Big\}=0$. Taking into account [7, Theorem 8.3], we can assert that $\psi(x/2)<3/4\psi(x)$ for a sufficient large integer $x$, thus
@@
In consequence, taking into account that $ax<\psi(x)<cx$ for $a=0.780355$... and $c=1.170533$..., for a sufficiently large $x$ (see [7, Theorem 8.3]), we deduce that
@@
Step 3. Next we deal with a sum over primes (see [7, Theorem 8.8(b)]).
@@
Let us obtain two coarse bounds. We know that the second term is bounded and, since $\log p/p^{n}$ is positive, it increases when the $x$ grows and the partial sums in $n$ grows. Moreover it is bounded from above by $\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}$ (see proof of [7, Theorem 8.8(b)]), which is bounded by the integral of $\dfrac{\log(x-1)}{(x-1)(x-2)}$ between $2$ and infinity. So, for sufficiently large $x$
@@
(6) $1.645<O_{3}(1)<-0.048<0.585-0.633$
Step 4. Finally we study the error term of the sum of reciprocal of primes (see [7, Theorem 8.8(d)]). We define $R(x)=\displaystyle\sum_{p\leq x}\dfrac{\log p}{p}-\log x$, so thanks to the estimate given in previous steps we get $R(x)=O_{3}(1)$. Let
@@
(7) $O(1/\log x)=\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\,.$
@@
Remark 2.3.
Estimates above are trivial, but we have seen that $R(x)$ is decreasing almost everywhere (a.e.) and negative for sufficiently large $x$. Let the expression of $b$ given in proof of [3, Theorem 427] (see also [2, Propostion]), that is
$\begin{split}b&=\int_{2}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt+1-\log\log 2\\ &=\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt+1-\log\log 2+\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\,.\end{split}$
Applying Stieltjes integration by parts a.e., we obtain
$\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt=\frac{R(2)}{\log 2}-\frac{R(x)}{\log x}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt=-1-\frac{R(x)}{\log x}+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,$,
Moreover $\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt$ is negative and strictly decreasing that converges to $b+\log\log 2-1$, hence
$\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt>b+\log\log 2-1\,$,
for sufficiently large $x$. Notice that $R(x)$ is derivable and integrable (a.e.) and so is its derivative, moreover $R(2)/\log 2=-1/2$. Therefore
$\begin{split}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt&=-b+1-\log\log 2+\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt\\ &=-b+1-\log\log 2-1-\frac{R(x)}{\log x}+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,.\end{split}$
With these things in hand we see that
$\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt=-b-\log\log 2+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,.$
However, from these equations we just can deduce that $\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}=\log\log 2-\log\log x$.
Remark 2.4.
Assume that
$O(1/\log x)=\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt=\frac{C(x)}{\log x}\,$,
thus
$\begin{split}C(x)&=\log x\cdot O(1/\log x)=R(x)-\log x\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\\ &=\sum_{p\leq x}\frac{\log p}{p}-\log x\Big(1+\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\Big)\end{split}$
and $\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt$ goes to zero when $x$ grows to infinity. In consequence $C(x)$ goes to $R(x)$ that is negative and bounded from above by a negative number (see (6)), that is $C(x)<C$ for some $C<0$